# Intuitionistic NF set theory

### v5

@@ 10. Results about ${\mathbb{T}}$
Proof. Suppose $m\in{\mathbb{F}}$. Then $\exists x\,(x\in 2^{{\mathbb{T}}m})$, by Lemma 10.14. Then by the definition of exponentiation, for some $u$ we have
$SSC(u)\in 2^{{\mathbb{T}}m}\ \land\ USC(u)\in{\mathbb{T}}m.$
@@
$USC(u)\in{\mathbb{T}}m$   by definition of ${\mathbb{T}}$
$SSC(u)\in{\sf FINITE}$   by Lemma 3.17
$SSC(u)\in 2^{{\mathbb{T}}m}$   by Lemma 8.12
@@
Proof. Suppose $m\in{\mathbb{F}}$. Then
@@
Proof. Suppose $2^{m}$ is inhabited. Then there exists $a$ with $USC(a)\in m$. Then
$2^{m}=Nc(SSC(a))$   by Lemma 7.2
$SSC(a)\in 2^{m}$   by Lemma 8.12
$USC(SSC(a))\in{\mathbb{T}}(2^{m})$   by Lemma 10.2
$USC(USC(a))\in{\mathbb{T}}m$   by Lemma 10.2
@@
$Nc(SSC(USC(a)))=Nc(USC(SSC(a)))$   by Lemma 6.2
$2^{Tm}=Nc(SSC(USC(a)))$   by Lemma 10.4
${\mathbb{T}}(2^{m})=Nc(USC(SSCa))$   by Lemma 10.3
@@
Proof. Suppose ${\mathbb{T}}n={\mathbb{T}}m$. By Lemma 4.7, we can find $a\in n$ and $b\in m$. Then
$USC(a)\in{\mathbb{T}}n$   by definition of ${\mathbb{T}}$
$USC(b)\in{\mathbb{T}}m$   by definition of ${\mathbb{T}}$
@@
$USC(a)\in{\mathbb{T}}n$   by the previous two lines
@@
Proof. The formula is stratified, giving $a$, $b$, and $c$ all index zero. Therefore we may proceed by induction on $b$.
@@
$({\mathbb{T}}a+{\mathbb{T}}b)^{+}\in{\mathbb{F}}$   by Lemmas  10.8 and 8.2
@@
That completes the induction step. That completes the proof of the lemma.
@@
$n<m{\ \leftrightarrow\ }{\mathbb{T}}n<{\mathbb{T}}m.$
Proof. Left to right is Lemma 10.12.
@@
Case 3, $m<n$. Then ${\mathbb{T}}m<{\mathbb{T}}n$ by Lemma 10.12. That completes the proof of the lemma.
@@
Proof. By induction on $p$ we will prove
(71) $\forall q\in{\mathbb{F}}\,(p<{\mathbb{T}}q\ {\rightarrow}\ \exists r\in{\mathbb{F}}\,(p={\mathbb{T}}r))$
@@
That completes the induction step. That completes the proof of the lemma.
@@
Proof. Suppose $p\in{\mathbb{F}}$ and $2^{p}$ is inhabited. Then by definition of exponentiation, for some $a$ we have $USC(a)\in p$ and $SSC(a)\in 2^{p}$. By definition of ${\mathbb{T}}$ we have $p={\mathbb{T}}(Nc(a))$. By Lemma 4.21, we have $Nc(a)\in{\mathbb{F}}$. That completes the proof of the lemma.
@@
Proof. By Theorem 5.18, $e={\sf zero}\ \lor\ e\neq{\sf zero}$. If $e={\sf zero}$ then $e^{+}={\sf one}$, so we are done by Lemma 4.20. Therefore we may assume $e\neq{\sf zero}$. By Lemma 4.7, $e+e$ is inhabited. By the definition of $<$, there exist $x$ and $y$ with $x\in e$ and $y\in e$ and $x\cap y=\Lambda$. Then
$y\neq\Lambda$   since if $y=\Lambda$ then $e={\sf zero}$, by Lemma 5.7
@@
Proof. The formula is stratified, giving $b$ and $c$ index 0 and $a$ index 1. ${\mathbb{F}}$ is just a parameter, so it does not need an index. Therefore we can proceed by induction on $a$.
@@
That completes the induction step. That completes the proof of the lemma.
@@
Proof. By definition, ${\sf zero}=\{\Lambda\}$. By the definition of addition, there exist $a$ and $b$ with $a\in p$ and $b\in q$ and $a\cap b=\Lambda$, such that $a\cup b\in{\sf zero}$. Then $a\cup b=\Lambda$. It follows that $a=\Lambda$ and $b=\Lambda$. On the other hand, if $a$ or $b$ had a non-empty member, then by the definition of addition, $a+b$ would have a non-empty member, so ${\sf zero}$ would have a non-empty member. Therefore $p=q=\{\Lambda\}={\sf zero}$. That completes the proof of the lemma.
@@
For $x,y\in{\mathbb{F}}$,
$x+x=y+y\ {\rightarrow}\ x=y.$
Proof. The formula is stratified; we prove it by induction on $x$, in the form
@@
That completes the induction step. That completes the proof of the lemma.
@@
Proof. Suppose $p\in{\mathbb{F}}$. Left to right:
@@
$USC(a)\in p$   for some $a$, by the definition of exponentiation
$USC(a)\in{\sf FINITE}$   by Lemma 4.4
@@
$Nc(a)\in{\mathbb{F}}$   by Lemma 4.21
$a\in Nc(a)$   by Lemma 4.11
$USC(a)\in{\mathbb{T}}(Nc(a))$   by definition of ${\mathbb{T}}$
${\mathbb{T}}(Nc(a))\in{\mathbb{F}}$   by Lemma 10.6
$USC(a)\in p\cap{\mathbb{T}}(Nc(a))$   by definition of intersection
$p={\mathbb{T}}(Nc(a))$   by Lemma 5.7
$\exists q\in{\mathbb{F}}\,(p={\mathbb{T}}a)$   namely $q=Nc(a)$
@@
$USC(u)\in{\mathbb{T}}q$   by definition of ${\mathbb{T}}$
@@
$SSC(u)\in 2^{{\mathbb{T}}q}$   by definition of exponentiation
$SSC(u)\in 2^{p}$   since $p={\mathbb{T}}q$
@@
That completes the proof of the right-to-left direction. That completes the proof of the lemma.
11. Cartesian products
The Cartesian product of two sets is defined as usual; the definition is stratified, so it can be given in INF. But because ordered pairs raise the types by two, the cardinality of $A\times B$ is not the product of the cardinalities of $A$ and $B$, but instead it is the product of ${\mathbb{T}}^{2}$ of those cardinalities. In this section we provide a proof of this fact, in the interest of setting down the fundamental facts about the theory of finite sets.

### v6

@@ 10 Results about $\mathbb{T}$}
Proof.
Suppose $m\in{\mathbb{F}}$. Then $\exists x\,(x\in 2^{{\mathbb{T}}m})$, by Lemma 10.14. Then by the definition of exponentiation, for some $u$ we have
$\mathcal{P}_{s}(u)\in 2^{{\mathbb{T}}m}\ \land\ \mathcal{P}_{1}(u)\in{\mathbb{T}}m.$
@@
$\mathcal{P}_{1}(u)\in{\mathbb{T}}m$   by definition of $\mathbb{T}$}
$\mathcal{P}_{s}(u)\in{\sf FINITE}$   by Lemma 3.17
$\mathcal{P}_{s}(u)\in 2^{{\mathbb{T}}m}$   by Lemma 8.12
@@
Proof.
Suppose $m\in{\mathbb{F}}$. Then
@@
Proof.
Suppose $2^{m}$ is inhabited. Then there exists $a$ with $\mathcal{P}_{1}(a)\in m$. Then
$2^{m}=|\mathcal{P}_{s}(a|)$   by Lemma 7.2
$\mathcal{P}_{s}(a)\in 2^{m}$   by Lemma 8.12
$\mathcal{P}_{1}(\mathcal{P}_{s}(a))\in{\mathbb{T}}(2^{m})$   by Lemma 10.2
$\mathcal{P}_{1}(\mathcal{P}_{1}(a))\in{\mathbb{T}}m$   by Lemma 10.2
@@
$|\mathcal{P}_{s}(\mathcal{P}_{1}(a|))=|\mathcal{P}_{1}(\mathcal{P}_{s}(a|))$   by Lemma 6.2
$2^{Tm}=|\mathcal{P}_{s}(\mathcal{P}_{1}(a|))$   by Lemma 10.4
${\mathbb{T}}(2^{m})=|\mathcal{P}_{1}(\mathcal{P}_{s}a|)$   by Lemma 10.3
@@
Proof.
Suppose ${\mathbb{T}}n={\mathbb{T}}m$. By Lemma 4.7, we can find $a\in n$ and $b\in m$. Then
$\mathcal{P}_{1}(a)\in{\mathbb{T}}n$   by definition of $\mathbb{T}$}
$\mathcal{P}_{1}(b)\in{\mathbb{T}}m$   by definition of $\mathbb{T}$}
@@
$\mathcal{P}_{1}(a)\in{\mathbb{T}}n$   by the previous two lines
@@
Proof.
The formula is stratified, giving $a$, $b$, and $c$ all index zero. Therefore we may proceed by induction on $b$.
@@
$({\mathbb{T}}a+{\mathbb{T}}b)^{+}\in{\mathbb{F}}$   by Lemmas 10.8 and 8.2
@@
That completes the induction step. ∎
@@
[…]
Proof.
Left to right is Lemma 10.12.
@@
Case 3, $m<n$. Then ${\mathbb{T}}m<{\mathbb{T}}n$ by Lemma 10.12. ∎
@@
Proof.
By induction on $p$ we will prove
$\forall q\in{\mathbb{F}}\,(p<{\mathbb{T}}q\ {\rightarrow}\ \exists r\in{\mathbb{F}}\,(p={\mathbb{T}}r))$ (73)
@@
That completes the induction step. ∎
@@
Proof.
Suppose $p\in{\mathbb{F}}$ and $2^{p}$ is inhabited. Then by definition of exponentiation, for some $a$ we have $\mathcal{P}_{1}(a)\in p$ and $\mathcal{P}_{s}(a)\in 2^{p}$. By definition of $\mathbb{T}$} we have $p={\mathbb{T}}(|a|)$. By Lemma 4.21, we have $|a|\in{\mathbb{F}}$. ∎
Lemma 10.23.
For $n,m\in{\mathbb{F}}$, we have
$n\leq m{\ \leftrightarrow\ }{\mathbb{T}}n\leq{\mathbb{T}}m.$
Proof.
We have
$n\leq m{\ \leftrightarrow\ }n<m\ \lor\ n=m$   by Lemma 5.19
${\mathbb{T}}n\leq{\mathbb{T}}m{\ \leftrightarrow\ }{\mathbb{T}}n<{\mathbb{T}}m\ \lor\ {\mathbb{T}}n={\mathbb{T}}m$   by Lemma 5.19
Now to prove the desired conclusion:
Left to right : if $n<m$ then ${\mathbb{T}}n<{\mathbb{T}}m$ by Lemma 10.20, so ${\mathbb{T}}n\leq{\mathbb{T}}m$. And if $n=m$, then ${\mathbb{T}}n={\mathbb{T}}m\leq{\mathbb{T}}m$, by Lemma 5.18.
Right to left : if ${\mathbb{T}}n<{\mathbb{T}}m$ then $n<m$ by Lemma 10.20, so $n\leq m$. And if ${\mathbb{T}}n={\mathbb{T}}m$, then $n=m$ by Lemma 10.18. ∎
@@
Proof.
By Theorem 5.17, $e={\sf zero}\ \lor\ e\neq{\sf zero}$. If $e={\sf zero}$ then $e^{+}={\sf one}$, so we are done by Lemma 4.20. Therefore we may assume $e\neq{\sf zero}$. By Lemma 4.7, $e+e$ is inhabited. By the definition of <<, there exist $x$ and $y$ with $x\in e$ and $y\in e$ and $x\cap y=\emptyset$. Then
$y\neq\emptyset$   since if $y=\emptyset$ then $e={\sf zero}$, by Lemma 4.24
@@
Proof.
The formula is stratified, giving $b$ and $c$ index 0 and $a$ index 1. $\mathbb{F}$} is just a parameter, so it does not need an index. Therefore we can proceed by induction on $a$.
@@
That completes the induction step. ∎
Lemma 10.26 (Specker 5.4).
Let $m\in{\sf NC}$. Then
$m\neq{\mathbb{T}}(m)+{\sf one}$
Remark. And so on, with $\sf one$} replaced by $\sf two$} or 23, 457, and any number you could name. If ${\mathbb{T}}m\neq m$, ${\mathbb{T}}m$ must be a non-standard distance away from $m$.
Proof.
We give the proof for $\sf one$}. Recall that $m$ is even if $m=p+p$ for some $p\in{\mathbb{F}}$, and odd if $m=p+p+{\sf one}$ for some $p\in{\mathbb{F}}$. Then if $m$ is even, $m+{\sf one}$ is odd, and vice versa. One can verify by induction that every integer is either or odd, and not both. Suppose ${\bf m}={\mathbb{T}}{\bf m}+{\sf one}$. If $m$ is even, then ${\mathbb{T}}m$ is even, by Lemma 10.25, so ${\mathbb{T}}(m)+{\sf one}$ is odd, contradiction. If $m$ is odd, then $m=k^{+}=k+{\sf one}$ for some $k\in{\mathbb{F}}$, since $\sf zero$} is even. Then $k$ is even. Then ${\mathbb{T}}(m)={\mathbb{T}}(k^{+})={\mathbb{T}}(k+{\sf one})={\mathbb{T}}(k)+{\sf one}$, which is odd since ${\mathbb{T}}(k)$ is even. Then ${\mathbb{T}}(m)+{\sf one}$ is even, contradiction, since $m$ is odd and equal to ${\mathbb{T}}(m)+{\sf one}$. ∎
@@
Proof.
By definition, $\sf zero$$\{\emptyset\}$. By the definition of addition, there exist $a$ and $b$ with $a\in p$ and $b\in q$ and $a\cap b=\emptyset$, such that $a\cup b\in{\sf zero}$. Then $a\cup b=\emptyset$. It follows that $a=\emptyset$ and $b=\emptyset$. On the other hand, if $a$ or $b$ had a non-empty member, then by the definition of addition, $a+b$ would have a non-empty member, so $\sf zero$} would have a non-empty member. Therefore $p=q=\{\emptyset\}={\sf zero}$. ∎
@@
For $\in{\mathbb{F}}, x+x=y+y$x=yx+x=y+y\ $\rightarrow$}\ x=y.
Proof.
The formula is stratified; we prove it by induction on $x$, in the form
@@
That completes the induction step. ∎
@@
Proof.
Suppose $p\in{\mathbb{F}}$. Left to right:
@@
$\mathcal{P}_{1}(a)\in p$   for some $a$, by the definition of exponentiation
$\mathcal{P}_{1}(a)\in{\sf FINITE}$   by Lemma 4.4
@@
$|a|\in{\mathbb{F}}$   by Lemma 4.21
$a\in|a|$   by Lemma 4.11
$\mathcal{P}_{1}(a)\in{\mathbb{T}}(|a|)$   by definition of $\mathbb{T}$}
${\mathbb{T}}(|a|)\in{\mathbb{F}}$   by Lemma 10.6
$\mathcal{P}_{1}(a)\in p\cap{\mathbb{T}}(|a|)$   by definition of intersection
$p={\mathbb{T}}(|a|)$   by Lemma 4.24
$\exists q\in{\mathbb{F}}\,(p={\mathbb{T}}a)$   namely $q=|a|$
@@
$\mathcal{P}_{1}(u)\in{\mathbb{T}}q$   by definition of $\mathbb{T}$}
@@
$\mathcal{P}_{s}(u)\in 2^{{\mathbb{T}}q}$   by definition of exponentiation
$\mathcal{P}_{s}(u)\in 2^{p}$   since $p={\mathbb{T}}q$
@@
That completes the proof of the right-to-left direction. ∎
Definition 10.30.
Let $X$ be any set of cardinals. Then we define
${\mathbb{T}}``(X)=\{{\mathbb{T}}(u):u\in X\}.$
or more explicitly
${\mathbb{T}}``(X)=\{{\mathbb{T}}(u):u\in X\}=\{y:\exists u\in X\,(y={\mathbb{T}}u)\}.$
The formula in the definition is stratified, giving $u$ index 0 and $y$ and $X$ index 1. Actually, $X$ is just a parameter and does not even need an index. Therefore the definition is legal in $\mathit{i}\mkern 0.4mu\mathrm{NF}$. We note that it is not a function definable in $\mathit{i}\mkern 0.4mu\mathrm{NF}$. It is just an abbreviation for a comprehension term. Note also that the set $X$ can be finite or not, and the cardinals in $X$ can be finite or not.
In general images commute with union. For images under $\mathbb{T}$} we have
Lemma 10.31.
${\mathbb{T}}``(X\cup Y)={\mathbb{T}}``(X)\cup{\mathbb{T}}``(Y)$.
Proof.
This is proved in a few short steps from the definitions of T(X)T``(X) and $\cup$. ∎
Lemma 10.32.
Let $a$ and $b$ be finite disjoint sets. Then
$|a\cup b|=|a|+|b|.$
Proof.
Left to right: Suppose $t\in|a\cup b|$. Then
$a\cup b\in|a\cup b|$   by Lemma 4.11
$a\in|a|$   by Lemma 4.11
$b\in|b|$   by Lemma 4.11
$a\cap b=\emptyset$   by hypothesis
$a\cup b\in{\sf FINITE}$   by Lemma 3.11
$|a|\in{\mathbb{F}}$   by Lemma 4.21
$|b|\in{\mathbb{F}}$   by Lemma 4.21
$|a\cup b|\in{\mathbb{F}}$   by Lemma 4.21
$|a|+|b|\in{\mathbb{F}}$   by Lemma 8.6
$a\cup b\in|a|+|b|$   by the definition of addition
$|a\cup b|=|a|+|b|$   by Lemma 4.24
Lemma 10.33.
Let $X$ be a finite set of cardinal numbers. Then
$|{\mathbb{T}}``(X|)={\mathbb{T}}(|X|).$
Proof.
The displayed formula in the lemma is stratified, giving $X$ index 1; then $|X|$ gets index 2 and ${\mathbb{T}}|X|$ gets index 3. On the left, the members of ${\mathbb{T}}``(X)$ are ${\mathbb{T}}u$ for $u\in X$, so $u$ gets index 0, and ${\mathbb{T}}u$ gets index 1, so ${\mathbb{T}}``X$ gets index 2, so $|{\mathbb{T}}``(X|)$ gets index 3, the same as the right side of the equation. So it is stratified, as claimed.
The part of the lemma involving $X$ is
$\forall X\,(X\in{\sf FINITE}\ {\rightarrow}\ X\subset NC\ {\rightarrow}\ \ {\rightarrow}\ |{\mathbb{T}}``(X|)={\mathbb{T}}(|X|),$
and this is also stratified, since $\sf FINITE$} and $NC$ are just parameters. Therefore we can prove it by induction on finite sets.
[…]
Induction step. Suppose $X$ is finite and $c\not\in X$. We have to show
$|{\mathbb{T}}``(X\cup\{c\}|={\mathbb{T}}(|X\cup\{c\}|.$
We have
$X\cup\{c\}\in{\sf FINITE}$ by Lemma 3.7, since $c\not\in X$
${\mathbb{T}}``(X\cup\{c\})={\mathbb{T}}``(X)\cup\{{\mathbb{T}}(c)\}$ by Lemma 10.31 (75)
$|{\mathbb{T}}``(X\cup\{c\}|)=|{\mathbb{T}}``(X|\cup\{{\mathbb{T}}(c)\})$ by the preceding line
${\mathbb{T}}c\not\in{\mathbb{T}}``(X)$ by Lemma 10.18, since $c\not\in X$ (76)
$|{\mathbb{T}}``(X|)={\mathbb{T}}(|X|)$ by the induction hypothesis (77)
$|X|\in{\mathbb{F}}$   by Lemma 4.21, since $X\in{\sf FINITE}$
${\mathbb{T}}(|X|)\in{\mathbb{F}}$   by Lemma 10.6
${\mathbb{T}}``(X)\in{\sf FINITE}$   by Lemma 4.4
$\{{\mathbb{T}}(c)\}\in{\sf FINITE}$   by Lemma 3.9
${\mathbb{T}}``(C)\cap\{T(c)\}=\emptyset$   by (76)
$|{\mathbb{T}}``(X|\cup\{{\mathbb{T}}(c)\}))=|{\mathbb{T}}``(X|)+|\{{\mathbb{T}}(c|\}))$   by Lemma 10.32
$|{\mathbb{T}}``(X|\cup\{{\mathbb{T}}(c)\}))=|{\mathbb{T}}``(X|)+{\sf one}$   by Lemma 10.7
$|{\mathbb{T}}``(X|\cup\{{\mathbb{T}}(c)\}))={\mathbb{T}}(|X|)+{\sf one}$   by the induction hypothesis (77)
$|{\mathbb{T}}``(X|\cup\{{\mathbb{T}}(c)\}))={\mathbb{T}}(|X|)+{\mathbb{T}}({\sf one})$   since ${\mathbb{T}}({\sf one})={\sf one}$
$|X|+{\sf one}=|X\cup\{c\}|$   by Lemma 10.32 since $c\not\in X$
$|X|+{\sf one}\in{\mathbb{F}}$   by Lemma 8.6
$|{\mathbb{T}}``(X)\cup\{{\mathbb{T}}(c)\})|={\mathbb{T}}(|X|+{\sf one})$   by Lemma 10.13
$|{\mathbb{T}}``(X)\cup\{{\mathbb{T}}(c)\}|={\mathbb{T}}(|X\cup\{c\}|)$   by the preceding lines
$|{\mathbb{T}}``(X\cup\{c\})|={\mathbb{T}}(|X\cup\{c\}|)$   by (75)
That completes the induction step. ∎
Lemma 10.34.
Let $X$ be a finite set of cardinals. Then ${\mathbb{T}}``(X)$ is finite.
Proof.
Let $X$ be a finite set of cardinals. Then
$X\subseteq{\sf FINITE}$   by hypothesis
${\mathbb{T}}``(X)\in{\sf FINITE}$   by hypothesis
$|{\mathbb{T}}``(X|)={\mathbb{T}}(|X|)$   by Lemma 10.33 (78)
$|X|)\in{\mathbb{F}}$   by Lemma 4.21
${\mathbb{T}}(|X|)\in{\mathbb{F}}$   by Lemma 10.6 (79)
$|{\mathbb{T}}``(X|\in{\mathbb{F}}$   by (78) and (79)
${\mathbb{T}}``(X)\in|{\mathbb{T}}``(X|)$   by Lemma 4.11
${\mathbb{T}}``(X)\in{\sf FINITE}$   by Lemma 4.4
11 Cartesian products
The Cartesian product of two sets is defined as usual; the definition is stratified, so it can be given in $\mathit{i}\mkern 0.4mu\mathrm{NF}$. But because ordered pairs raise the types by two, the cardinality of $A\times B$ is not the product of the cardinalities of $A$ and $B$, but instead it is the product of ${\mathbb{T}}^{2}$ of those cardinalities. In this section we provide a proof of this fact, in the interest of setting down the fundamental facts about the theory of finite sets.